# Data-driven Quantum Dynamical Embedding Method for Long-term Prediction on Near-term Quantum Computers

### v3

@@ II.4 Experimental settings
The quantum circuit of the QDE comprises three primary components: encoding, parametrized evolution, and quantum measurement, as detailed in Sec. II.1. To simplify the QDE circuit, we implement the encoding layer $U_{enc}(\cdot)$ with rotation $R^{i}_{Y}(\arccos(\cdot))$ for the $i^{\text{th}}$ qubit. Here, $R_{Y}(\cdot)$ is defined as $R_{Y}(\theta)=e^{-i\frac{\theta}{2}\sigma_{y}}$, with the input normalized to the range $[-1,1]^{\otimes n}$. Furthermore, the observables $\mathcal{H}_{m},\mathcal{H}_{x}$ are designed as $\sigma_{z}$ for the respective qubit. This design choice is driven by the understanding that the input can be efficiently reconstructed using the $\sigma_{z}$ expectation, eliminating the need for additional evolution. As the number of qubits, circuit depth, and linear-composite channels increase, the QDE exhibits more complex dynamics. The methodology for scaling these elements is depicted in Fig. 1(c) and properties are further discussed in Sec. III.4. Throughout the time sequence iterations, all quantum circuit parameters remain constant.
@@
where $h$ represents the transverse field strength and $J_{ij}$ denotes the coupling strength of site $i$ and $j$. The coefficients $D_{i},J_{ij}$ are randomly chosen from a uniform distribution within $[-W,W]$ and $[-J_{s},J_{s}]$, respectively. And all remain fixed during training.
HEA [52]: This architecture utilizes a CZ entanglement layer combined with single-qubit rotations $U_{1}$ as described in TIEA. In the specific case of a 2-qubit QDE, the architecture simplifies to a CZ gate followed by single-layer parametrized $R_{Y}$ rotations, which are then utilized in the prediction of composite signals.
@@ III Results
Refer to captionRefer to caption(a)(b)(c)(d)(e)(f)
Figure 2: Demonstrations of QDE for Time Series Prediction: (a) 2-qubit QDE application on cosine-wave prediction with $x(t)=0.5\cos(\omega t)$, where $\omega=\pi/25$ and the value of time axis is rescaled by $\Delta=0.04$. The training and predicting lengths are set to $L=T=100$ in this and the following tests. (b) Composite periodic signal predicting, $x(t)=0.2\cos(\omega t)+0.3\sin(2\omega t)$, where $\omega=\pi/25$. (c) Aperiodic time series predicting, $x(t)=0.2\cos(\omega t)+0.3\sin(\sqrt{5}\omega t)$, with the same architecture and settings as (b). (d), (e), and (f) correspond to the noisy versions of the scenarios described in (a), (b), and (c), respectively, trained with noise uniformly distributed within $[-0.1,0.1]$. For these versions, the number of training points is increased to enhance the capture of signal characteristics buried in noise. The prediction part of each subplot is juxtaposed with theoretical clean signals for comparison. The red dashed line represents the deviation $\delta=\hat{x}(t)-x_{\text{ref}}(t)$, where $\hat{x}(t)$ is the output of the QDE model and $x_{\text{ref}}(t)$ is the reference, the noisy signal in training stage and theoretical signal in predicting stage.
@@ III.1 Applications
Figure 3: Comparison of QDE with QRC on Cosine-Wave Signal Prediction. (a) Error map for different values of the Hamiltonian hyper-parameters of QRC, with the transverse field $h$ and the disorder bound $W$ varying logarithmically in the range $[10^{-2},10^{2}]$. Results are averaged over 100 realizations. (b) and (c) With $h/J_{s}$ ($W/J_{s}$) fixed at the minimum error from map (a), the other parameter $W/J_{s}$ ($h/J_{s}$) is varied. The average results are shown with the minimum to maximum error range (grey shadows) and standard deviation (red shadows). The QDE model’s performance is consistently below the average error and standard deviation of the QRC model.
This part establishes both numerical and experimental evidence on some applications to characterize the adaptability of the QDE introduced above. Firstly, we employ the QDE to predict a cosine-wave signal with a single channel. In the context of composite signal functions, our approach effectively captures the characteristics inherent in superpositions of different cosine-wave signals with the architecture depicted in Fig. 1(c). To further demonstrate the QDE’s versatility, we assess its denoising capability on signals exposed to uniform noise. For benchmarking purposes, we compare it with the performance using QRC. Additionally, we focus on long-term time series and noisy simulations, training and forecasting them at a ratio of $1:10$ under the influence of depolarizing or amplitude-damping noise. Notably, an experimental test on a superconducting quantum computer indicates significant promise for real-world time series processing.
We employ the HEA architecture in this application and tailor the number of linear-composite channels to align with the task’s complexity.
Cosine-wave signal
@@
where $\omega=0.04\pi$, $A=0.5$ and $\phi=0$ for simplicity. In this application, the QDE configuration is simplified to a single channel comprising two qubits. We define the temporal domain as $0\leq t\leq 200,t\in\mathbb{N}$. In this and the following demonstrations, we train the QDE with the first 100 points generated from the equation (light red dots) and test the extrapolation prediction result for the next 100 points (blue triangles), as depicted in Fig. 2(a). The prediction error, defined by the MSE, is $1.10\times 10^{-5}$, which demonstrates the capability of QDE to accurately represent the periodic nature of the cosine-wave. Detailed theoretical foundations of these observations are further elaborated in Sec. III.4.
Composite signals
We consider applying the QDE on the composite cosine-waves
@@
where each term represents an independent wave component. In Fig. 2(b) and (c), we demonstrate two cases, a periodic and an aperiodic one with only two linear-composite channels, which can be viewed as two simplest non-trivial examples to benchmark the model. In the first case, we set $\omega_{1}=0.04\pi$ and $\omega_{2}=2\omega_{1}$; the second case, $\omega_{1}=0.04\pi$ and $\omega_{2}=\sqrt{5}\omega_{1}$. The corresponding prediction errors are $5.21\times 10^{-6}$ and $8.10\times 10^{-5}$, respectively. In both cases, each channel of the QDE follows the ansatz configuration used in cosine-wave prediction. The decent performance on prediction error proved that the QDE correctly learned the systems’ dynamics.
@@
To evaluate the denoising capability of the QDE, we expose the aforementioned three signals to uniform noise randomly distributed within the range of $[-0.1,0.1]$. Each noisy signal undergoes processing via the QDE to evaluate its efficacy in recovering the original, clean signal. The number of training points in these versions is increased to enhance the capture of signal characteristic mode buried in random noise. For the cosine-wave, the QDE markedly reduces noise interference, achieving a denoising MSE of $3.07\times 10^{-3}$, as illustrated in Fig. 2(d). Likewise, the periodic and aperiodic composite signals demonstrate the QDE’s robustness in noise reduction, recording MSEs of $3.97\times 10^{-3}$ and $3.21\times 10^{-3}$, respectively, as shown in Fig. 2(e) and (f). The prediction part of each subplot is juxtaposed with theoretical clean signals for comparison. We define the deviation as $\delta=\hat{x}(t)-x_{\text{ref}}(t)$, which represents the difference between the output of the QDE ($\hat{x}(t)$) and the reference signal value ($\hat{x}(t)$). Although $\delta$ is random and large during the training stage, as indicated by the red dashed line, the QDE effectively extracts the noise-free mode from the signal, showcasing its notable performance in denoising. These results not only underscore the QDE’s effectiveness in signal fidelity restoration but also highlight its utility in practical noise reduction scenarios.
@@
In this application, we compare the performance of the QDE and QRC models on the aforementioned tasks. For the QRC model, we utilize a 3-qubit system with 5 evolution slices ($V=5$), keeping the training and prediction lengths consistent with those of the QDE model. The results of this comparison for the first task, cosine-wave signal prediction, are shown in Fig. 3(a)-(c). The error map for different values of the Hamiltonian hyper-parameters of QRC is plotted with the transverse field $h$ and the disorder bound $W$ varying logarithmically in the range $[10^{-2},10^{2}]$. Results are averaged over 100 realizations. Fig. 3(b) and Fig. 3(c) demonstrate that when $h/J_{s}$ ($W/J_{s}$) is fixed at the minimum error from the map in Fig. 3(a) and the other parameter $W/J_{s}$ ($h/J_{s}$) varies in the range $[10^{-2},10^{2}]$, the MSE of the QDE model, indicated by the dashed light blue line, consistently remains below the average error (red curve) and standard deviation (red shadows) of the QRC model. Despite some instances of lower prediction error across 100 realizations, the 2-qubit QDE model exhibits more stable performance in terms of precision, circuit depth, and training resource allocation compared to the QRC model. Further results on the other two tasks, along with detailed insights into the QRC theory and its configurations, can be found in the supplementary information.
@@ III.2 Performing long-term prediction by recovering from noisy results
We evaluated long-term prediction capabilities of the QDE under aforementioned noise channels using a cosine-wave signal and two composite signals. Maintaining a training-to-predicting ratio of $1:10$ and replicating previous settings, including the LECL, these simulations highlight the robustness of the QDE complemented by a linear layer. This robustness demonstrates the QDE’s potential for long-term time series prediction in practical scenarios.
@@
Figure 5: Error Distribution in Encoding Space and the Cosine-wave Signal Generation. (a) Error distribution within the encoding space $(m,x)$, each grid point indicating the deviation between the ideal output $(m^{*},x^{*})$ and the quantum computer’s output $(\hat{m},\hat{x})$ without LECL, denoted as $\epsilon_{m}=\hat{m}-m^{*}$ and $\epsilon_{x}=\hat{x}-x^{*}$. (b) Adjusted error distribution following the application of LECL, illustrating the changes in error magnitude. (c) Generation of a single period of a cosine-wave time series, showing the data register dynamics and the memory register responses; solid lines correspond to the results from ideal simulations with optimized parameters, while dashed lines represent the outcomes from the quantum computer after the LECL is applied.

### v4

@@ II.4 Experimental settings
The quantum circuit of the QDE comprises three primary components: encoding, parametrized evolution, and quantum measurement, as detailed in Sec. II.1. To simplify the QDE circuit, we implement the encoding layer $U_{enc}(\cdot)$ with rotation $R^{i}_{Y}(\arccos(\cdot))$ for the $i^{\text{th}}$ qubit. The input is normalized to the range $1,1]^{\otimes n}$. This design choice is driven by the understanding that the input can be efficiently reconstructed using the $\sigma_{z}$ expectation, eliminating the need for additional evolution. As the number of qubits, circuit depth, and linear-composite channels increase, the QDE exhibits more complex dynamics. The methodology for scaling these elements is depicted in Fig. 1(c) and properties are further discussed in Sec. III.4. Throughout the time sequence iterations, all quantum circuit parameters remain constant.
@@
where $h$ represents the transverse field strength and $J_{ij}$ denotes the coupling strength of site $i$ and $j$. The coefficients $D_{i},J_{ij}$ are randomly chosen from a uniform distribution within [W,W][-W,W] and $J_{s},J_{s}]$, respectively. And all remain fixed during training.
HEA [59]: This architecture can be viewed as a variant of TIEA, where the Hamiltonian evolution layer is replaced with a CZ entanglement layer, while retaining the single-qubit rotation gates $U_{1}$. For certain time series tasks discussed in the next section, the rotation gate can be further simplified. For certain time series tasks discussed in the next section, the rotation gate can be further simplified.
@@ III Results
Refer to captionRefer to captionRefer to caption(a)(b)(c)(d)(e)(f)(g)
Figure 2: Demonstrations of QDE for Time Series Prediction: (a) $(1,1)$-QDE application on cosine-wave prediction with $x(t)=0.5\cos(\omega t)$, where $\omega=\pi/25$ and the value of time axis is rescaled by $\Delta=0.04$. The training and predicting lengths are set to $L=T=100$ in this and the following tests, unless otherwise specified. (b) Composite periodic signal predicting, $x(t)=0.2\cos(\omega t)+0.3\sin(2\omega t)$, where $\omega=\pi/25$. (c) Aperiodic time series predicting, $x(t)=0.2\cos(\omega t)+0.3\sin(\sqrt{5}\omega t)$, with the same architecture and settings as (b). (d), (e), and (f) correspond to the noisy versions of the scenarios described in (a), (b), and (c), respectively, trained with noise uniformly distributed within [0.1,0.1][-0.1,0.1]. For these versions, the number of training points is increased to enhance the capture of signal characteristics buried in noise. The prediction part of each subplot is juxtaposed with theoretical clean signals for comparison. The red dashed line represents the deviation $\delta=\hat{x}(t)-x_{\text{ref}}(t)$, where $\hat{x}(t)$ is the output of the QDE model and $x_{\text{ref}}(t)$ is the reference, the noisy signal in training stage and theoretical signal in predicting stage. (g) Prediction of NARMA2 data using the $(2,1)$-QDE model. The training length is set to 500 steps, with a fixed prediction length. For clarity, only the last 200 training points are shown. The signal is generated following the formula described in the main text.
@@ III.1 Applications
Refer to caption(a)(b)(c)(d)(e)
Figure 3: Performance Scaling of the QDE Model for Cosine-wave Signal Prediction Under Two Scenarios. (a)-(c) The first row demonstrates the effects of increasing the total number of channels in a $(1,1)$-QDE configuration. (d)-(e) The second row illustrates performance changes when scaling up the number of internal qubits within a single QDE block (i.e., increasing $n_{x}$ and $n_{m}$). Each violin plot represents the distribution of MSE outcomes across 50 trials, highlighting median trends and variability associated with different circuit depths and qubit counts.
Models Tasks Qubits Channels Depth Gates Parameters MSE
QDE Cosine-wave 2 1 1 5 2 $1.10\times 10^{-6}$
Composite-periodic signal 4 2 1 10 4 $5.21\times 10^{-6}$
Composite-aperiodic signal 4 2 1 10 4 $8.10\times 10^{-5}$
QRC All the above 3       26 See Fig. 4 (S5 and S6)
Table 1: Quantum Resources Used in QDE and QRC Models. The QRC model is implemented through Hamiltonian evolution $e^{-iHt}$ rather than an explicit gate-based circuit. Therefore, resource metrics used in the QDE model such as channel count, single-block depth, and gate count are not applicable to QRC and are indicated as “ ”. In the QDE model, comparisons are made by scaling the number of single $(1,1)$-QDE blocks, referred to as channels. The ansatz employed is the HEA as described in Sec. II.4.
@@
Figure 4: Comparison of QDE with QRC on Cosine-Wave Signal Prediction. (a) Error map for different values of the Hamiltonian hyper-parameters of QRC, with the transverse field $h$ and the disorder bound $W$ varying logarithmically in the range $10^{-2},10^{2}]$. Results are averaged over 100 realizations. (b) and (c) With $h/J_{s}$ ($W/J_{s}$) fixed at the minimum error from map (a), the other hyper-parameter $W/J_{s}$ ($h/J_{s}$) is varied. The average results are shown with the minimum to maximum error range (grey shadows) and standard deviation (red shadows). The reference line for the QDE model represents the average performance over different circuit parameter initializations using a single $(1,1)$-QDE channel. Since the QDE used here is based on a HEA architecture without hyper-parameters, it is naturally independent of the x-axis values.
Figure 5: Performance of the QDE model on NARMA2. The four subfigures are arranged from left to right by increasing qubit count. Each violin plot represents the distribution over 50 trials, consistent with previous tests.
This part establishes both numerical and experimental evidence on some applications to characterize the adaptability of the QDE introduced above. Firstly, we employ the QDE to predict a cosine-wave signal with a single channel. In the context of composite signal functions, our approach effectively captures the characteristics inherent in superpositions of different cosine-wave signals with the architecture depicted in Fig. 1(c). To further demonstrate the QDE’s versatility, we assess its denoising capability on signals exposed to uniform noise. For benchmarking purposes, we compare it with the performance using QRC. Due to the existence of local minima in the QDE optimization landscape, we report the median performance over multiple initial parameter sets. The distribution of these performances is illustrated using a violin plot, as shown in Fig. 3, taking the cosine-wave signal as an example. In contrast, the QRC model, trained via linear regression, demonstrates less sensitivity to initial parameters; thus, we report the average performance over 100 realizations varying only hyper-parameters, excluding training parameters. Additionally, we focus on long-term time series and noisy simulations, training and forecasting them at a ratio of $1:10$ under the influence of depolarizing or amplitude-damping noise. Notably, an experimental test on a superconducting quantum computer indicates significant promise for real-world time series processing.
We employ the HEA architecture in these applications  and tailor the number of linear-composite channels to align with each task’s complexity. By default, to evaluate per-step predictive accuracy and enable fair comparisons, we report one-step predictions, i.e., at each time step the ground-truth value is provided as input, consistent with standard QRC evaluations. Nevertheless, because QDE is designed for multi-step forecasting, we also provide autoregressive (multi-step) simulation results and analyses in Secs. III.2III.4. For datasets in which the input and target signals arise from different feature sets (e.g., NARMA), we report only one-step prediction results throughout this work.
Cosine-wave and its composite
@@
where $\omega=0.04\pi$, $A=0.5$ and $\phi=0$ for simplicity. Then we consider the composite of Cosine-waves,
@@
where each term represents an independent wave component. Two simplest non-trivial examples will be presented, a periodic and an aperiodic one with only two linear-composite channels. In the periodic case, we set $\omega_{1}=0.04\pi$ and $\omega_{2}=2\omega_{1}$; the aperiodic case, $\omega_{1}=0.04\pi$ and $\omega_{2}=\sqrt{5}\omega_{1}$. We define the temporal domain as $0\leq t\leq 200,t\in\mathbb{N}$. In this and the following demonstrations, we train the QDE using the first 100 data points generated from the equation (light red dots) and evaluate its extrapolation performance on the subsequent 100 points (blue triangles), as shown in Fig. 2. The ansatz used here is a simplified version of HEA, in which the three-parameter gate $U_{1}(\theta_{1},\theta_{2},\theta_{3})$ is replaced with a single-parameter rotation $R_{Y}(\theta)$ to investigate the minimal quantum resources required for signal prediction. We begin by identifying the lowest resource configuration for the three signals, as summarized in Table I: one $(1,1)$-QDE channel is sufficient for the cosine wave, while two channels are used for the other two signals. The prediction error, defined by the MSE and averaged on 50 samples with different initializations, is $1.10\times 10^{-6}$, $5.21\times 10^{-6}$ and $8.10\times 10^{-5}$ with the lowest resource, respectively, which demonstrates the capability of QDE to accurately represent the periodic nature of each component of the signal. Detailed theoretical foundations of these observations are further elaborated in Sec. III.4.
Then, using the cosine-wave prediction as an example, we analyze the scalability of the QDE model by either increasing the total number of $(1,1)$-QDE channels or increasing the number of qubits within a single $(n_{m},n_{x})$-QDE block. The first method combines multiple independent $(1,1)$-QDE outputs classically without entanglement, while the second increases internal qubit count to enable entanglement within the block, creating a more complex quantum state. Due to local minima in optimization, violin plots are used to display the full MSE distribution. Results in Fig. 3 show that increasing the channel number significantly reduces MSE, whereas the performance gains from increasing layers eventually saturate, dependent on the channel number. Similarly, increasing internal qubits improves best-case performance but with greater variability and instability.
@@
To evaluate the denoising capability of the QDE, we expose the aforementioned three signals to uniform noise randomly distributed within the range of [0.1,0.1][-0.1,0.1]. Each noisy signal undergoes processing via the QDE to evaluate its efficacy in recovering the original, clean signal. The number of training points in these versions is increased to enhance the capture of signal characteristic mode buried in random noise. For the cosine-wave, the QDE markedly reduces noise interference, achieving a denoising MSE of $3.07\times 10^{-3}$, as illustrated in Fig. 2(d). Likewise, the periodic and aperiodic composite signals demonstrate the QDE’s robustness in noise reduction, recording MSEs of $3.97\times 10^{-3}$ and $3.21\times 10^{-3}$, respectively, as shown in Fig. 2(e) and (f). The prediction part of each subplot is juxtaposed with theoretical clean signals for comparison. We define the deviation as $\delta=\hat{x}(t)-x_{\text{ref}}(t)$, which represents the difference between the output of the QDE ($\hat{x}(t)$) and the reference signal value ($\hat{x}(t)$). Although $\delta$ is random and large during the training stage, as indicated by the red dashed line, the QDE effectively extracts the noise-free mode from the signal, showcasing its notable performance in denoising. These results not only underscore the QDE’s effectiveness in signal fidelity restoration but also highlight its utility in practical noise reduction scenarios.
@@
In this application, we compare the performance of the QDE and QRC models on the aforementioned tasks. For the QRC model, we utilize a 3-qubit system with 5 evolution slices ($V=5$), keeping the training and prediction lengths consistent with those of the QDE model. The results of this comparison for the first task, cosine-wave signal prediction, are shown in Fig. 4(a)-(c). The error map for different values of the Hamiltonian hyper-parameters of QRC is plotted with the transverse field $h$ and the disorder bound $W$ varying logarithmically in the range $10^{-2},10^{2}]$. Results are averaged over 100 realizations. In Fig. 4(b)-(c), we fix $h/J_{s}$ (or $W/J_{s}$) at the optimal value identified in Fig. 4(a). The other parameter ($W/J_{s}$ or $h/J_{s}$) is varied within the range $10^{-2},10^{2}]$. The median MSE of the QDE model is calculated from 50 random initializations, as detailed in Fig. 3 and is indicated by the dashed light blue line. The median MSE consistently remains below the QRC model’s average error (red curve) and standard deviation (red shaded area). Detailed configuration parameters and resource calculations for both models are summarized in Table 1, representing the minimal resources required to accomplish the task. Despite some instances of lower prediction error across 100 realizations, the 2-qubit QDE model exhibits more stable performance in terms of precision, circuit depth, and training resource allocation compared to the QRC model. Further results on the other two tasks, along with detailed insights into the QRC theory and its configurations, can be found in the supplementary information.
NARMA2
To further demonstrate the capability of the QDE model on more complex time series, we apply it to the nonlinear autoregressive moving average (NARMA) task. A typical NARMA signal of order $n$ is generated by the following dynamics:
$x_{t+1}=...\alpha y_{t}+\beta y_{t}\left(\sum_{j=0}^{n-1}y_{t-j}\right)+\gamma s_{t-n+1}s_{t}+\delta,$
where $x_{t}$ denotes $x(t)$ for simplicity, and the parameters $(\alpha,\beta,\gamma,\delta)$ are set to $(0.3,0.05,1.5,0.1)$. The input signal $s_{t}$ is superimposed sine wave, following the setup in [22]:
$s_{t}=0.1\left[\sin\left(\frac{2\pi at}{T}\right)\sin\left(\frac{2\pi bt}{T}\right)\sin\left(\frac{2\pi ct}{T}\right)+1\right],$
with $(a,b,c)=(2.11,3.73,4.11)$ and $T=100$. In this simulation, we slightly modify the input and output protocol, as shown in Fig. 1, to process the input $(\bm{m}_{t},s_{t})$ to output $(\bm{m}_{t+1},x_{t+1})$, aligning with the settings in [22]. We set $n=2$, train for 500 steps, and predict for 100 steps. Given the complexity of this time series, we chose to scale the model by increasing the number of internal qubits, rather than simply adding more channels, and by deepening the ansatz layers to fully exploit entanglement among qubits, aiming to improve model expressiveness. All results are over 50 trails with different initializations. Results using the ansatz HEA for NARMA2 are shown in Fig. 5, with a demonstration of a $(2,1)$-QDE provided in Fig. 2(g). We observe that increasing the number of ansatz layers consistently reduces the error, while increasing qubit count does not always guarantee better performance due to increased optimization complexity. This example highlights the QDE model’s potential for handling more realistic and challenging time series beyond simpler composite signals. For other complex tasks like learning and predicting the nonlinear dynamics of the Rayleigh equation, please refer to the supplementary information.
The Problem of Barren Plateau
As discussed in Sec. II.3, the potential issue of barren plateaus in the QDE model is an important consideration. To investigate this, we use the NARMA2 signal as a test case and compute both the variance of the loss function and the variance of its gradients. The circuit setups are consistent with those described previously. The results, shown in Fig. 6, present these variances averaged over 400 samples with parameters $\theta_{i}\in[0,2\pi)$. From Fig. 6, we observe that when the number of ansatz layers is small, the exponential decay of variance with increasing qubit number is not apparent. However, as the number of layers increases, this decay trend becomes more evident, and at 20 layers, the data closely follow an exponential decay pattern consistent with barren plateau behavior. This indicates that the QDE model is susceptible to barren plateaus when the ansatz depth is large. Although the layer depths used in our current experiments are not sufficient to fully enter the barren plateau regime, the phenomenon remains a limitation of the model. Addressing methods to mitigate or avoid barren plateaus is beyond the scope of this study and will be the focus of our future research efforts.
Refer to captionRefer to caption(a)(b)
Figure 6: Barren Plateau Evaluation of the $(n_{m},1)$-QDE Model Using the HEA The total qubit number is scaled from 2 to 8. The variance of the loss function $\langle\mathcal{L}(\bm{\theta})\rangle$ (a) and its gradient $\langle\partial_{\mu}\mathcal{L}(\bm{\theta})\rangle$ for specified index $\mu$ (b), are shown varying with qubit number for different ansatz layer depths and plotted with a semi-log style. Each data point represents an average over 400 samples of parameters drawn uniformly from $0,2\pi)$.
@@ III.2 Performing long-term prediction by recovering from noisy results
We evaluated long-term prediction capabilities of the QDE under aforementioned noise channels using composite Cosine-wave signals. To demonstrate long-term forecasting performance and robustness to accumulated errors, here we employ a multi-step (autoregressive) prediction protocol for composite signals: each newly predicted value is fed back as input for the next time step. The aperiodic signal is shown in Fig. 7, with additional examples provided in the supplementary material. Maintaining a training-to-predicting ratio of $1:10$ and replicating previous settings, including the LECL, these simulations highlight the robustness of the QDE complemented by a linear layer. This robustness demonstrates the QDE’s potential for long-term time series prediction in practical scenarios.
@@
[…]